OpenStudy (anonymous):

if,cos root y=x,then show (1-x^2)d^2y/dx^2-x dy/dx-2=0 ????????

OpenStudy (anonymous):

\[\cos(\sqrt{y})=x\] right? i guess you have to find \[\frac{dy}{dx}\] and \[\frac{d^2y}{dx^2}\] and check. is there supposed to be a snap way to do it?

OpenStudy (anonymous):

i've got ittt thanks